# A nonperturbative test of nucleation calculations for strong phase transitions

### v2

@@ I Introduction
First order phase transitions in the early universe have been of sustained interest. A primordial phase transition could help to explain the matter-antimatter asymmetry of the universe [1, 2, 3, 4, 5, 6]. The phase transition itself, as well as any resulting nonequilibrium physics of the primordial plasma, would produce a stochastic gravitational wave background that could potentially be detected [7].
Hints of a gravitational wave background have been seen by the various Pulsar Timing Array (PTA) collaborations [8, 9, 10, 11]. Such a gravitational wave background is expected to be due to supermassive black holes [12], but it is not possible to rule out other new physics such as a phase transition [13]. Future missions such as LISA will be well placed to look for gravitational waves from electroweak-scale phase transitions [14, 15, 16], but the theoretical uncertainties in the predicted gravitational wave power spectrum will need to be constrained.
Thanks in part to large-scale simulations [17, 18], considerable progress has been made in modelling the resulting gravitational wave power spectrum [19, 20]. At the same time, the accuracy of the quantities which parametrise the power spectrum – including the nucleation rate – have faced growing scrutiny [21, 22, 23, 24, 25, 26].
Experimental tests of nucleation theory in condensed matter systems show a mixed picture. For the AB transition in superfluid 3He, there is a longstanding and puzzling discrepancy [27, 28, 29], whereas good agreement was found for nucleation in a ferromagnetic superfluid [30]. Recently, there have been proposals to test bubble nucleation in ultracold atomic gases [31, 32, 33, 34, 35, 36, 37].
There is a need for reliable predictions with controlled systematic uncertainties so that we can test particle physics models against gravitational wave observations, as well as for comparison with analogue experiments. In this paper we focus on calibrating the accuracy of nucleation rate calculations in relativistic finite-temperature field theory, comparing perturbative semiclassical methods against direct nonperturbative numerical simulation.
@@
In the field theoretical context – at zero temperature – a metastable vacuum state can decay via quantum tunnelling [38]. A complete one-loop calculation of the rate of this process was first carried out by Coleman and Callan [39, 40], using a saddlepoint approximation of the path integral. This result has since been rederived from a number of perspectives and formalisms [41, 42, 43].
At high temperature, bubbles may also nucleate classically. Within classical field theory, a complete one-loop calculation for this process was first carried out by Langer [44, 45, 46], extending the tree-level theory of Cahn and Hilliard [47]. However, the analogous calculation within quantum field theory is much less clear. Early papers by Linde [48] and Affleck [49] gave slightly different expressions in the high temperature regime. Neither agrees with the classical result of Langer.
The above methods all depend on a semiclassical picture of the bubble and its fluctuations. A fully numerical, non-perturbative lattice calculation of the bubble nucleation rate was introduced in Ref. [50]. At the time, the focus was on the physics of the minimal Standard Model, where a first-order phase transition would have arisen from radiative corrections to the quartic Higgs potential. In Ref. [51] a toy model with similar features was studied. More recently, in Ref. [52] a more complete study of nucleation in the minimal Standard Model was carried out, motivated by the idea that any sufficiently heavy new particles could be integrated out [22].
Another approach has been to study thermal nucleation directly as a real time process [53, 54, 55, 56, 57, 58, 59, 60]. This involves evolving the lattice in time, waiting until a growing bubble appears and recording the time taken for the bubble to appear. Since the nucleation rate can vary over many orders of magnitude, this approach is only viable over a narrow temperature range.
@@
We will focus on the dynamics of a single bubble and its real time evolution from the metastable phase to the stable phase. The bubbles arise as long wavelength thermal fluctuations of the scalar field. These fluctuations are highly occupied in a thermal bath, and so their effective evolution is classical [65, 66, 67, 68]. The effects of short wavelength fluctuations appear both through screening of the effective parameters, and as noise and damping in the evolution equations [69, 68, 70].
@@
The effective Hamiltonian, $H_{\text{eff}}$, is constructed to match the long-wavelength equal-time correlation functions of the full quantum theory [71, 72]. This matching reduces the long-wavelength equilibrium thermodynamics of a four-dimensional (4d) quantum field theory to that of a 3d classical statistical field theory, and hence is known as high-temperature dimensional reduction (see Section I in the Supplemental Material for details). Importantly, for a scalar field theory, an appropriate choice of $\gamma$ also leads to the matching of the unequal time correlation functions [65, 66, 67, 68].
@@ II Saddlepoint approximation
For field theories, the computation of this probability current was described by Langer [44, 45, 46]. The relationship of Langer’s formalism to high-temperature quantum field theories has been studied in Refs. [75, 76], and generalised to higher orders in Ref. [77]. A crucial result is that the nucleation rate factorises
@@
where $A_{\text{dyn}}$ is the dynamical factor and $A_{\text{stat}}$ is the statistical factor. The latter is a purely time-independent quantity, equal to the vacuum nucleation rate in 2+1 dimensions. At leading (one-loop) order, $2\pi A_{\text{dyn}}$ is equal to the initial growth rate of the critical bubble. Note that while this split into dynamical and statistical parts mirrors the split in our lattice approach, there are nevertheless differences.
@@
However, for applications of nucleation theory, especially for more complicated models, neither the functional determinant nor the dynamical prefactor are typically computed. In Section II in the Supplemental Material we show two common lower-order approximations: in the first, $\Gamma_{\text{tree-level}}$, the functional determinant and dynamical prefactor are estimated as $T^{4}$, and in the second, $\Gamma_{\text{LPA}}$, a local potential approximation is made for the functional determinant, together with a prescription for how to drop unwanted imaginary parts.
@@ III Lattice simulations
Given the limitations of directly simulating the Hamiltonian system of Eqs. (3) and (4), our nonperturbative calculation of the nucleation rate closely follows the approach introduced in Refs. [50, 51].
@@
where the sum over $x$ extends over the 3d lattice sites. We have removed the cubic term by a constant shift in the field. One can derive exact lattice-continuum relations for this model within lattice perturbation theory [82, 83], and improvements up to $O(a^{2})$ [84, 51, 85]. See Section III in the Supplemental Material for more information about our lattice action.
@@ III.1 Critical bubble probability
The separatrix configurations are suppressed by $\sim e^{-65}$ at the temperature we simulate. Our lattice Monte Carlo simulations therefore use the multicanonical method [86, 87] to overcome this suppression and generate a sufficient number of near-critical bubble configurations (see Ref. [61] for its application to the current model). The order parameter $\theta_{\mathrm{op}}$ is measured, yielding a histogram that (below the critical temperature) will resemble Fig. 1. If the choice of order parameter, system volume and geometry admit, then there will be a local minimum between the metastable and stable phase peaks that we identify with the separatrix. From this local minimum, an estimate of $\theta_{\mathrm{op}}$ corresponding to critical bubble configurations, $\theta_{\mathrm{c}}$, can be inferred. Configurations sufficiently near the separatrix $\theta_{\mathrm{op}}\in[\theta_{\mathrm{c}}-\frac{\epsilon}{2},\theta_{\mathrm{c}}+\frac{\epsilon}{2}]$ are used as the initial conditions for evolution of the the lattice versions of the stochastic evolution equations (3) and (4), along with momenta drawn from a Gaussian distribution. The final result is independent of the exact choice of $\epsilon$, but it should be sufficiently small that the sampled configurations correspond to near-critical bubbles.
@@ III.3 Results
Given our use of an $O(a^{2})$ improved lattice discretisation, the leading dependence on lattice spacing arises at $O(a^{3})$. For a fixed volume of $L\lambda_{3}=42$ we vary the lattice spacing and perform a least-squares fit $f(a)=b+ca^{3}$ to the logarithm of the rate. The best fit line, with $\chi^{2}/\mathrm{dof}=5.53$, is plotted in the figure. Jackknife resampling this fit yields $\log(\Gamma/\lambda_{3}^{4})=-73.26(10)$, where the error is dominated by our largest lattice spacing $a\lambda_{3}=3.5$. This is approximately the inverse screening mass of the system, at which point the cubic form is expected to break down. A one-loop estimate of the screening mass yields $m_{\text{s}}^{\text{PT}}/\lambda_{3}=0.294(1)$ [61]. From the figure we deduce that lattice artefacts are comparable to statistical uncertainties already with $\lambda_{3}a=1.5$.
For reference we also show the nucleation rate computed using the linear order parameter $\theta_{\mathrm{op}}=\phi_{\text{lin}}$ for one lattice spacing. We find good agreement (within errors) with our final results, however our final choice of order parameter shows substantially reduced errors, which we attribute to the suppression of the metastable phase bulk fluctuations by our new choice of order parameter.
We then show the dependence on volume at constant lattice spacing $\lambda_{3}a=1.5$. Away from the second-order phase transition, this 3d Euclidean model has no massless modes, so long-range correlations die off exponentially with distance. Motivated by this, we fit the logarithm of the rate to an exponential $f(L)=b+c\exp(-m_{\text{s}}L)$, finding $m_{\text{s}}/\lambda_{3}=0.287(12)$, consistent with the perturbative screening mass. We show the best fit line in the figure, with $\chi^{2}/\mathrm{dof}=1.54$. Jackknife resampling this fit yields $\log(\Gamma/\lambda_{3}^{4})=-74.08(5)$ for the infinite-volume extrapolation. Note that fitting $f(L)=b+c/L^{n}$ with $n\in\{1,2,3\}$ to our finite-volume data yields much poorer fits with $\chi^{2}/\text{d.o.f.}\sim 10^{2}$.
@@
Figure 3: The zero lattice spacing (top) and infinite volume (bottom) extrapolations, together with cubic and exponential fits respectively. The vertical line in the top plot marks a perturbative estimate of the correlation length.
Figure 4: The nucleation rate as a function of temperature. Uncertainty bands for the tree-level and one-loop perturbative results are based on varying the renormalisation scale over $\mu_{3}/\lambda_{3}\in\{0.5,1,2\}$. The uncertainty estimate on the LPA relflects different choices for removing imaginary parts of the potential. The lattice points are continuum extrapolated; the red triangle highlights the temperature actually simulated, while the orange circles have utilised reweighting. The orange continuous line is the reweighted result for $a\lambda_{3}=1.5$, $L\lambda_{3}=60$. The results in this figure are tabulated at https://doi.org/10.5281/zenodo.10891523.
@@ IV Discussion
The main result of this work is the logarithm of the nucleation rate, reliably computed to high precision for a real scalar quantum field theory at high temperature. Statistical uncertainties are much smaller than discrepancies with other methods, as seen in Fig. 4. Systematic uncertainties related to the continuum extrapolation are also well under control, as evidenced by Fig. 3. This has been made possible by use of efficient multicanonical algorithms [86, 87, 50, 51], and exact and $O(a^{2})$ improved lattice-continuum relations [82, 83, 84, 51, 85].
The benchmark parameter point that we simulated was chosen with the hope of testing perturbation theory in a regime where it is expected to work well. At the critical temperature, the tree-level approximation to the discontinuity in the background field value agrees with the lattice to 10% accuracy, and the one-loop approximation agrees to better than 1% [61].
Yet, for $|\log\Gamma|$ at the temperature simulated, we have found that the lattice result is 100% larger than the tree-level approximation and almost 20% larger than the one-loop approximation. It may be that this is explained by the growth of the loop expansion parameter as the system supercools, which becomes almost twice as large at the temperature we have simulated than at the critical temperature. Note that the discrepancy in $|\log\Gamma|$ cannot be explained by a shift of the critical temperature between orders, as in this model the critical temperature is fixed by the $Z_{2}$ symmetry between phases, and hence is exact already at tree-level within the 3d EFT [61].
@@
While our lattice simulations give reliable and precise predictions for the nucleation rate, their calculation is slow. To use such lattice simulations for phenomenological studies of cosmological phase transitions, faster sampling of the relevant bubble configurations is needed. We have identified one direction of progress, through the development of order parameters optimised to enhance the sampling of tunnelling configurations. Our quadratic order parameter $\theta_{\mathrm{op}}=\bar{\phi}^{2}-2A\bar{\phi}$ was crucial for efficiently simulating the larger volumes. In a subsequent paper we will explore this further and also set out how our computations can be applied to predictions of the gravitational wave spectrum.
@@ Acknowledgements.
The authors would like to thank Kari Rummukainen for his insights and advice at a number of points throughout this study. O.G. (ORCID ID $0$-0002-7815-3379) was supported by the Research Funds of the University of Helsinki, U.K. Science and Technology Facilities Council (STFC) Consolidated Grant ST/T000732/1, a Research Leadership Award from the Leverhulme Trust, and a Royal Society Dorothy Hodgkin Fellowship. A.K. (ORCID ID $0$-0002-0309-3471) was supported by Research Council of Finland grant no. 328958 and a travel grant from the Jenny and Antti Wihuri Foundation. D.J.W. (ORCID ID $0$-0001-6986-0517) was supported by Research Council of Finland grant nos. 324882, 328958, 349865 and 353131. The authors also wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.

### v3

@@ I Introduction
First order phase transitions in the early universe have been of sustained interest. A primordial phase transition could help to explain the matter-antimatter asymmetry of the universe [1, 2, 3]. The phase transition itself, as well as any resulting nonequilibrium physics of the primordial plasma, would produce a stochastic gravitational wave background that could potentially be detected [4].
Hints of a gravitational wave background have been seen by the various Pulsar Timing Array (PTA) collaborations [5, 6, 7, 8]. Such a gravitational wave background is expected to be due to supermassive black holes [9], but it is not possible to rule out other new physics such as a phase transition [10]. Future missions such as LISA will be well placed to look for gravitational waves from electroweak-scale phase transitions [11, 12], but the theoretical uncertainties in the predicted gravitational wave power spectrum will need to be constrained.
Thanks to large-scale simulations [13, 14], considerable progress has been made in modelling the resulting gravitational wave power spectrum [15, 16]. At the same time, the accuracy of the quantities which parametrise the power spectrum – including the nucleation rate – have faced growing scrutiny [17, 18, 19, 20, 21, 22].
Experimental tests of nucleation theory in condensed matter systems show a mixed picture. For the AB transition in superfluid 3He, there is a longstanding and puzzling discrepancy [23, 24, 25], whereas good agreement was found for nucleation in a ferromagnetic superfluid [26]. Recently, there have been proposals to test bubble nucleation in ultracold atomic gases [27, 28, 29].
There is a need for reliable predictions with controlled systematic uncertainties so that we can test particle physics models against gravitational wave observations, as well as for comparison with analogue experiments. We focus on calibrating the accuracy of nucleation rate calculations in relativistic finite-temperature field theory, comparing perturbative semiclassical methods against direct nonperturbative numerical simulation.
@@
In field theory at zero temperature, a metastable vacuum state can decay via quantum tunnelling [30]. A complete one-loop calculation of the rate of this process was first carried out by Coleman and Callan [31, 32], using a saddlepoint approximation of the path integral. This result has since been rederived from a number of perspectives and formalisms [33, 34, 35].
At high temperature, bubbles can nucleate classically. Within classical field theory, a complete one-loop calculation for this process was first carried out by Langer [36, 37], extending the tree-level theory of Cahn and Hilliard [38]. However, the analogous calculation within quantum field theory is much less clear. Early papers by Linde [39] and Affleck [40] gave slightly different expressions in the high temperature regime. Neither agrees with the classical result of Langer.
The above methods all depend on a semiclassical picture of the bubble and its fluctuations. A fully numerical, lattice calculation of the bubble nucleation rate was introduced in Ref. [41]. At the time, the focus was on the physics of the minimal Standard Model, where a first-order phase transition would have arisen from radiative corrections to the quartic Higgs potential. In Ref. [42] a toy model with similar features was studied. Recently a more complete study of nucleation in the minimal Standard Model was carried out [43], motivated by the idea that any sufficiently heavy new particles could be integrated out [18].
Another approach has been to study thermal nucleation directly as a real time process [44, 45, 46, 47, 48, 49, 50]. This involves evolving the lattice in time, waiting until a growing bubble appears and recording the time taken for the bubble to appear. Since the nucleation rate can vary over many orders of magnitude, this approach is only viable over a narrow temperature range.
@@
We focus on the dynamics of a single bubble and its real time evolution from the metastable phase to the stable phase. The bubbles arise as long wavelength thermal fluctuations of the scalar field. These fluctuations are highly occupied in a thermal bath, and so their effective evolution is classical [55, 56, 57]. The effects of short wavelength fluctuations appear both through screening of the effective parameters, and as noise and damping in the evolution equations [58, 57].
@@
The effective Hamiltonian, $H_{\text{eff}}$, is constructed to match the long-wavelength equal-time correlation functions of the full quantum theory [59, 60]. This matching reduces the long-wavelength equilibrium thermodynamics of a four-dimensional (4d) quantum field theory to that of a 3d classical statistical field theory, and hence is known as high-temperature dimensional reduction (see Section I in the Supplemental Material for details). Importantly, for a scalar field theory, an appropriate choice of $\gamma$ also leads to the matching of the unequal time correlation functions [61, 55, 56].
@@ II Saddlepoint approximation
For field theories, the computation of this probability current was described by Langer [36, 37]. The relationship of Langer’s formalism to high-temperature quantum field theories has been studied in Refs. [64, 65], and generalised to higher orders in Ref. [66]. A crucial result is that the nucleation rate factorises
@@
where $A_{\text{dyn}}$ is the dynamical factor and $A_{\text{stat}}$ is the statistical factor. The latter is a purely time-independent quantity, equal to the vacuum nucleation rate in 2+1 dimensions. At leading (one-loop) order, $2\pi A_{\text{dyn}}$ is equal to the initial growth rate of the critical bubble.
@@
However, for applications of nucleation theory, especially for more complicated models, neither the functional determinant nor the dynamical prefactor are typically computed. In Section II in the Supplemental Material we show two common lower-order approximations: in the first, $\Gamma_{\text{tree-level}}$, the functional determinant and dynamical prefactor are estimated as $T^{4}$, and in the second, $\Gamma_{\text{LPA}}$, a local potential approximation is made for the functional determinant, together with a prescription for dropping unwanted imaginary parts.
@@ III Lattice simulations
Our nonperturbative calculation of the nucleation rate closely follows the approach introduced in Refs. [41, 42].
@@
where the sum over $x$ extends over the 3d lattice sites. We have removed the cubic term by a constant shift in the field. One can derive exact lattice-continuum relations for this model within lattice perturbation theory [71], and improvements up to $O(a^{2})$ [72, 42, 73]. See Section III in the Supplemental Material for more information about our lattice action and [74] for the simulation code.
@@ III.1 Critical bubble probability
The separatrix configurations are suppressed by $\sim e^{-65}$ at the temperature we simulate. Our Monte Carlo simulations therefore use the multicanonical method [75, 76] to overcome this suppression and generate a sufficient number of near-critical bubble configurations (see Ref. [51] for its application to the current model). The order parameter $\theta_{\mathrm{op}}$ is measured, yielding a histogram that (below $T_{c}$) will resemble Fig. 1. If the choice of order parameter, system volume and geometry admit, then there will be a local minimum between the metastable and stable phase peaks that we identify with the separatrix, at $\theta_{\mathrm{op}}=\theta_{\mathrm{c}}$. This corresponds to critical bubble configurations. Configurations sufficiently near the separatrix $\theta_{\mathrm{op}}\in[\theta_{\mathrm{c}}-\frac{\epsilon}{2},\theta_{\mathrm{c}}+\frac{\epsilon}{2}]$ are used as the initial conditions for evolution of the the lattice versions of the stochastic evolution equations (3) and (4), along with momenta drawn from a Gaussian distribution. The final result is independent of the exact choice of $\epsilon$, but it should be sufficiently small that the sampled configurations correspond to near-critical bubbles.
@@ III.3 Results
Given our use of an $O(a^{2})$ improved lattice discretisation, the leading dependence on lattice spacing arises at $O(a^{3})$. For a fixed volume of $L\lambda_{3}=42$ we vary the lattice spacing and perform a least-squares fit $f(a)=b+ca^{3}$ to the logarithm of the rate. The best fit line, with $\chi^{2}/\mathrm{dof}=0.79$, excluding $\lambda_{3}a=3.5$, is plotted in the figure. Jackknife resampling this fit yields $\log(\Gamma/\lambda_{3}^{4})=-73.24(11)$. Our largest lattice spacing $a\lambda_{3}=3.5$ is approximately the inverse screening mass of the system — a one-loop estimate of the screening mass yields $m_{\text{s}}^{\text{PT}}/\lambda_{3}=0.294(1)$ [51] — at which point the cubic form is expected to break down. Motivated by this we fit $f(a)=b+ca^{3}+da^{4}$ to all data points, giving $\chi^{2}/\mathrm{dof}=1.76$. We find agreement on the fit parameters between the two approaches. From the figure we deduce that lattice artefacts are comparable to statistical uncertainties already with $\lambda_{3}a=1.5$.
For reference we also show the nucleation rate computed using the linear order parameter $\theta_{\mathrm{op}}=\phi_{\text{lin}}$ for one lattice spacing. We find agreement within error with our final results, however our quadratic order parameter shows substantially reduced errors, which we attribute to the relative suppression of the metastable phase bulk fluctuations.
We then show the dependence on volume at constant lattice spacing $\lambda_{3}a=1.5$. Away from the second-order phase transition, this 3d Euclidean model has no massless modes, so long-range correlations die off exponentially with distance. Motivated by this, we fit the logarithm of the rate to an exponential $f(L)=b+c\exp(-m_{\text{s}}L)$, finding $m_{\text{s}}/\lambda_{3}=0.283(12)$, consistent with the perturbative screening mass. We show the best fit line in the figure, with $\chi^{2}/\mathrm{dof}=1.54$. Jackknife resampling this fit yields $\log(\Gamma/\lambda_{3}^{4})=-74.09(5)$ for the infinite-volume extrapolation. Note that fitting $f(L)=b+c/L^{n}$ with $n\in\{1,2,3\}$ to our finite-volume data yields much poorer fits with $\chi^{2}/\text{d.o.f.}\sim 10^{2}$.
@@
Figure 3: The zero lattice spacing (top) and infinite volume (bottom) extrapolations, together with cubic, cubic plus quartic and exponential fits respectively. The vertical line in the top plot marks a perturbative estimate of the correlation length.
Figure 4: The nucleation rate as a function of temperature. Uncertainty bands for the tree-level and one-loop perturbative results are based on varying the renormalisation scale over $\mu_{3}/\lambda_{3}\in\{0.5,1,2\}$. The uncertainty estimate on the LPA relflects different choices for removing imaginary parts of the potential. The lattice points are continuum extrapolated; the red triangle highlights the temperature actually simulated, while the orange circles have utilised reweighting. The orange continuous line is the reweighted result for $a\lambda_{3}=1.5$, $L\lambda_{3}=60$. The results in this figure are tabulated at [77].
@@ IV Discussion
The main result of this work is the logarithm of the nucleation rate, reliably computed to high precision for a real scalar quantum field theory at high temperature. Statistical uncertainties are much smaller than discrepancies with other methods, as seen in Fig. 4. Systematic uncertainties related to the continuum extrapolation are also well under control, as evidenced by Fig. 3. This has been made possible by use of efficient multicanonical algorithms [75, 76, 41, 42], and exact and $O(a^{2})$ improved lattice-continuum relations [71, 72, 42, 73].
The benchmark parameter point that we simulated was chosen with the hope of testing perturbation theory in a regime where it is expected to work well. [^1] For our chosen xSM parameter point in 4d, at one-loop we find $T_{\text{c}}\approx 98.5~\text{GeV}$, $\Delta\langle\varphi\rangle_{\text{c}}/T_{\text{c}}\approx 1.67$ and the percolation temperature is $T_{*}\approx 94.8~\text{GeV}$ [11]. The loop expansion converges rapidly for the field condensate (adopting units where $\lambda_{3}=1$[51]
$\Delta\langle\phi\rangle_{\text{c}}^{\text{lat}}=1.341(2),\quad\Delta\langle\phi\rangle_{\text{c}}^{\text{pt}}=1.2+0.1378+\dots.$ (16)
We have displayed the first two terms of the loop expansion within the 3d EFT. The tree-level prediction agrees with the lattice to 10%, and at one-loop to 0.2%. The two- and three-loop terms are smaller still. In Sec. III.3, we have also found that the lattice and one-loop estimates of the screening mass agree within statistical uncertainties.
Yet for the nucleation rate at the temperature simulated, and at the same physical parameter point, we have
$|\log\Gamma|^{\text{lat}}=74.09(5),\quad|\log\Gamma|^{\text{pt}}=38.0+25.3+\dots.$ (17)
One might expect that this lack of convergence could be explained by the growth of the loop expansion parameter as the system supercools, which becomes almost twice as large at the temperature we have simulated than at the critical temperature. However, even at twice larger couplings the discontinuity in the field condensate agrees to 20% at tree-level and 2% at one-loop [51]. Note that the discrepancy in the rate cannot be explained by an offset in temperature, as both lattice and perturbation theory use the same EFT parameters, and in this model the critical temperature is fixed by the $Z_{2}$ symmetry between phases, and hence is exact already at tree-level within the 3d EFT [51].
@@
While our lattice simulations give reliable and precise predictions for the nucleation rate, their calculation is slow. To use such lattice simulations for phenomenological studies of cosmological phase transitions, faster sampling of the relevant bubble configurations is needed. We have identified one direction of progress, through the development of order parameters optimised to enhance the sampling of tunnelling configurations. Our quadratic order parameter $\theta_{\mathrm{op}}=\bar{\phi}^{2}-2A\bar{\phi}$ was crucial for efficiently simulating the larger volumes.
@@ Acknowledgements.
The authors would like to thank Kari Rummukainen for his insights and advice at a number of points throughout this study. O.G. (ORCID ID 0000-0002-7815-3379) was supported by the Research Funds of the University of Helsinki, U.K. Science and Technology Facilities Council (STFC) Consolidated Grant ST/T000732/1, a Research Leadership Award from the Leverhulme Trust, and a Royal Society Dorothy Hodgkin Fellowship. A.K. (ORCID ID 0000-0002-0309-3471) was supported by Research Council of Finland grant no. 328958 and a travel grant from the Jenny and Antti Wihuri Foundation. D.J.W. (ORCID ID 0000-0001-6986-0517) was supported by Research Council of Finland grant nos. 324882, 328958, 349865 and 353131. The authors also wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.